If z1 = 2cis115° and z2 = 3cis65° , write z1z2 in its RECTANGULAR FORM: a + bi.

Answer by ikleyn(52781)

.
If z1 = 2cis115° and z2 = 3cis65° , write z1z2 in its RECTANGULAR FORM: a + bi.
~~~~~~~~~~~~~

z1 = 2cis(115°)  and  z2 = 3cis(65°),  then


    z1 * z2 = (2*3) * cis(115° + 65°) = 6*cis(180°) = -6 = -6 + 0*i.    ANSWER


So, in the form a + bi,  a = -6,  b = 0.    ANSWER

Solved, answered and carefully explained.